OpenStudy (anonymous):

Solve the system of equations using matrices. Use Gaussian elimination with back- substitution. x+y+z = -5 x-y+3z = -1 4x+y+z = -2

OpenStudy (anonymous):

x+y+z is wat

OpenStudy (anonymous):

@musically2332 what do you mean?

OpenStudy (anonymous):

oh sorry i was asking u its xyz

OpenStudy (anonymous):

okay so what does that mean?

OpenStudy (anonymous):

{x + y + z = -5 , x - y + 3 z = -1, 4 x + y + z = -2} ?

OpenStudy (anonymous):

@robtobey it an equation

OpenStudy (anonymous):

so how do i solve those equations @robtobey typed?

OpenStudy (anonymous):

you have to use like terms

OpenStudy (anonymous):

-5(x-y+3z 3zx-y

OpenStudy (anonymous):

-1 4x+y +z =-2

OpenStudy (anonymous):

okay now I'm really confused, what do i solve specifically